Press: \underline{q} to quit

Active-DVI

.PS
# The trigonometric circle
arrowhead=7
scale=1
d=0.3
d2=d/2
C: circle radius 3*d
"O" at C.w + (2.7*d,-0.3*d)
"The trigonometric circle" at C.s - (0, d)
I: "~1" ljust at C.e below
J: "1~" rjust at C.n left above
X: arrow from C.w - (0.5*d,0) to C.e + (1.5*d,0)
"$x$" rjust at X.e below
Y: arrow from C.s - (0,0.5*d) to C.n + (0,1.5*d)
"$y$~" rjust at Y.n
Orig : C.center
Theta: arc -> from Orig + (1.6*d,0) to Orig + (1.3*d,0.75*d)
"~$\theta$" ljust at Theta.e above
M :  Orig + (3*d*cos(0.5236),3*d*sin(0.5236))
line from Orig to M
"$M$" at M +(2*d2/3,2*d2/3)
line -> from M to (Orig.x,Orig.y+M.y)
"$\sin\left(\theta\right)\left\{%
  \begin{array}{c} \\ \\ \end{array}%
\right.$" at Orig + (-0.8*d,0.75*d)
line -> from M to (M.x,Orig.y)
"$\underbrace{\hspace*{1.95cm}}_%
 {\displaystyle\cos~\left(\theta\right)~}$" rjust
.PE
\showgraph